OpenStudy (anonymous):

Solve: 3(x + 4) − 4(x + 3) = 7x + 16

OpenStudy (vijay):

x= -2

OpenStudy (anonymous):

\[3x+12-4x-12=7x+16\] \[-x=7x+16\] \[-8x=16\] \[x = 16/-8\] \[x=-2\]

OpenStudy (anonymous):

Thankzx